(a) Find the value of Z C (x + y) ds, where C is parameterized by x = t, y = t and 0 ≤ t ≤ 1


1
Expert's answer
2021-11-03T17:01:18-0400
"\dfrac{dx}{dt}=1, \dfrac{dy}{dt}=1"

"\int_C(x+y)ds=\displaystyle\int_{0}^{1}(t+t)\sqrt{(1)^2+(1)^2}dt"

"=\sqrt{2}[t^2]\begin{matrix}
 1\\
 0
\end{matrix}=\sqrt{2}"

"\int_C(x+y)ds=\sqrt{2}"